# SpeechIQ: Speech Intelligence Quotient Across Cognitive Levels in Voice Understanding Large Language Models

### v1

@@ Abstract
SpeechIQ: Speech Intelligence Quotient Across Cognitive Levels in Voice Understanding Large Language Models
@@ 1 Introduction
The rapid rise of voice understanding Large Language Models (LLM) that can process voice input via the speech portal has ushered in a new paradigm of human-machine interaction Reddy (1988), where LLM process spoken instructions, infer semantic intent, and execute downstream tasks Jelinek et al. (1991); Jurafsky et al. (1995); Zue and Glass (2000); Tur et al. (2005); Mesnil et al. (2014); Kawahara (2019); Huang et al. (2024); Mahmood et al. (2025). LLM bridge speech and language intelligence Sparks et al. (1996), enabling applications from voice assistants to interactive robots. A fundamental prerequisite for deploying robust LLM lies in establishing reliable evaluation metrics for voice understanding. These metrics ensure that the system accurately comprehends voice inputs. Automatic speech recognition (ASR) is the central component of existing cascaded111We refer to two-pass based neural ASR-LM systems Hoffmeister et al. (2016); Sainath et al. (2019); Huang et al. (2020), which have been widely adopted in the current voice assistants, including Siri, Alexa, and Google Assistant. LLM systems, transcribing voice into text for subsequent LLM processing. Consequently, the research community primarily assesses voice understanding using transcription error metrics Hunt (1990); Klakow and Peters (2002), with Word Error Rate (WER) standing as the de facto standard.
@@
However, WER primarily measures the lexical recall of ASR, its transcribed quality does not fully capture the semantic understanding He et al. (2011); Chen et al. (2018); Szymański et al. (2020) and task completion Van Strien et al. (2020) capabilities of LLMs via their speech portal. For instance, WER treats all lexical errors equally, ignoring how transcription inaccuracies propagate to higher-level comprehension (i.e., semantic understanding) and executions (i.e., downstream tasks).
@@
Meanwhile, a growing trend in LLM research is the shift toward end-to-end multi-modal models Rubenstein et al. (2023); Radhakrishnan et al. (2023); Maiti et al. (2024); Hu et al. (2024); Défossez et al. (2024); Chen et al. (2024), which bypass explicit ASR transcription and directly map audio inputs to task-specific outputs (e.g., answers, actions). While this paradigm simplifies pipelines and reduces error cascades, its end-to-end voice service obviates the need for generating an intermediate transcript, rendering traditional WER-based evaluation obsolete Kuo et al. (2020); Tüske et al. (2021). This creates a critical methodological gap: without a reliable and potentially unified metric, comparing voice understanding across architectural choices, such as the modularity of “cascaded neural systems” and the efficiency of “end-to-end mutlimodal models,” remains insufficiently characterized.
@@
The above two limitations mirror a fundamental challenge in assessing artificial intelligence: as Figure 1, human-like intelligence is inherently hierarchical. Drawing on Bloom’s Taxonomy Bloom and Krathwohl (1969); Bengio et al. (2009), a cognitive framework where skills ascend from basic remembering to advanced creating, we posit that LLM should be evaluated through analogous levels. In this framework, WER represents the lowest level by measuring verbatim recall while neglecting higher-order abilities such as semantic understanding and task-solving. For instance, two transcripts with identical WER may differ drastically in meaning Szymański et al. (2023), leading to divergent LLM responses or failed instructions.
@@
To address this gap, we propose a hierarchical IQ test for LLM that evaluates intelligence across three fundamental levels (refer to limitations) aligned with Bloom’s Cognitive Taxonomy: (1) Remember in verbatim: Lexical accuracy is quantified using WER; (2) Understand: Semantic consistency is assessed by comparing LLM responses to ASR outputs and ground-truth transcripts. LLMs are prompted to infer the background context (e.g., domain, speaker intent) and generate a summary of the speech, then measure hidden states similarity between ASR-derived and ground truth-derived responses; (3) Apply: Task-solving capability is tested via multi-choice question-answering (QA) pairs constructed from ground-truth transcripts. LLM answers questions based on speech inputs, with accuracy reflecting real-world utility. Then, we draw inspiration from Raven’s Progressive Matrices Raven and Court (1998); John and Raven (2003) (i.e., a human IQ test) that aggregates performance across various dimensions into a single score.
@@ 2.1 ASR Evaluation
ASR systems have long been evaluated using lexical similarity metrics such as WER, which quantifies the Levenshtein distance Navarro (2001) between ASR outputs and reference transcripts. Metrics like character error rate MacKenzie and Soukoreff (2002) (CER), sentence error rate Juffs and Harrington (1996) (SER), and translation error rate (TER) extend this framework to the sentence or translation level but share WER’s core limitation: treating all errors equally, regardless of semantic impact. While effective for benchmarking transcription fidelity, these metrics do not fully capture how errors propagate to downstream tasks.
Recent work mitigates WER’s semantic insensitivity by incorporating sentence embeddings into evaluation. Sentence similarity-based metrics Kim et al. (2021, 2022) and BERTScore Zhang et al. (2020) leverage pre-trained language models to compute semantic correspondence between ASR hypotheses and ground-truth references. Others propose task-specific metrics, such as Medconcept WER Adedeji et al. (2024) which arranges more weights on medical entities in WER computation or severity score Whetten and Kennington (2023) that involves sentiment similarity in the evaluation. However, these methods are either confined to specific tasks or require curated labeled data, limiting their generalizability to open-domain LLM.
Efforts to develop hybrid metrics that combine error rate and semantic similarity, such as H_eval Sasindran et al. (2023), and Sema Sasindran et al. (2024), further enrich the evaluation landscape. Concurrently, machine translation inspired metrics like BLEU Score Papineni et al. (2002) and COMET Rei et al. (2020), have started to be adapted for ASR to assess fluency and pragmatic adequacy. However, these approaches still emphasize text-to-text alignment between ASR outputs and references, rather than from a view of influencing the responses and actions of downstream LLMs, motivating the proposed SpeechIQ as a multi-dimensional metric.
@@ 2.2 LLM Understanding Systems
In general, speech-based LLMs encompass both the understanding of voice signals and the generation Agostinelli et al. (2023); Borsos et al. (2023a); Yang et al. (2024c) of vocal or general audio outputs. In this work, we focus exclusively on voice understanding tasks with text outputs as one first step to examine voice intelligence Sparks et al. (1996) or its potential voice world model Ha and Schmidhuber (2018); Matsuo et al. (2022) via LLM backbone, which we denote as LLM. LLM understanding systems are designed to interact Shah et al. (2018); López et al. (2018) with users and execute task-oriented instructions via speech-based inputs. This form of “voice-to-text” architecture can be broadly categorized into three main types, all of which achieve state-of-the-art ASR quality on public leaderboards Radford et al. (2022); Chen et al. (2023); Puvvada et al. (2024):
@@
This approach follows a traditional pipeline-based structure, where an ASR model Watanabe et al. (2018); Gulati et al. (2020); Radford et al. (2022) transcribes the spoken input, and the 1-best transcription is passed to textual LLM as one operation system Wu et al. (2024); Dighe et al. (2024) for direct response generation.
@@
In this approach, the ASR model first generates multiple possible transcriptions (i.e., hypotheses) for a given speech input, where “texts” are then treated as input features for a many-to-one GER post-editing module Yang et al. (2023); Velikovich et al. (2024); Hori et al. (2025). The GER module refines these hypotheses to determine the most accurate transcription, which is subsequently processed by the LLM.
@@
These models are inherently designed to support both speech and text modalities Rubenstein et al. (2023); Hu et al. (2024); Lyu et al. (2023); Zhang et al. (2024a). Unlike cascaded approaches, they process audio inputs directly and generate textual outputs without requiring intermediate ASR transcriptions. This paradigm leverages a multimodal learning framework (i.e., a speech embedding encoder Radhakrishnan et al. (2023) or tokenizer Borsos et al. (2023b); Zhang et al. (2024b)) to achieve end-to-end speech understanding and reasoning.
In summary, we carefully selected and examined aforementioned three ASR-LLM systems. System 1 serves as the most cost-effective solution Hermansky and Junqua (1988) for voice interactions with LLMs. System 2 is designed to decode texts that offer richer information Chan et al. (2023); Yang et al. (2024b) for textual applications. System 3 highlights a low-latency, multimodal embedding injection approach for large language models OpenAI et al. (2024); Team et al. (2024). Next, we introduce our proposed method, which corresponds to Bloom’s cognitive taxonomy.
@@ 3.2 Understand: Similarity Metric
To quantify the impact of ASR errors on semantic comprehension, we measure the deviation in LLM responses caused by ASR transcription errors. Specifically, inspired by recent work (Jiang et al., 2024; Liu et al., 2024) showing that instructing LLMs to generate one word is a simple and straightforward approach to representing semantic meanings, we evaluate LLM-generated responses via two key questions: (1) : The background scenario of the speech [] in one word is. This helps resolve ambiguities and measure whether the ASR content preserves core contextual meaning; (2) : The summary of this speech [] in one word is. This assesses the extent to which ASR errors affect the LLM’s high-level understanding. Instead of directly comparing LLM-generated words, we use the last layer of hidden states of LLM for generating the next token as embeddings and compute their cosine similarity with the embeddings generated from the ground truth transcription:
@@ 3.3 Apply: QA Accuracy
During QA generation, we leverage GPT-4o and prompt it to focus on either the core concept or information details in the transcription. (Appendix A) To mitigate potential errors in generated QA, we also employ GPT-4o itself and Gemini-1.5-flash to answer these questions using the ground-truth transcription. Questions, where GPT-4o or Gemini-1.5-flash fails to produce correct answers, are discarded and regenerated. Then during evaluation, cascaded systems answer questions based on ASR transcriptions, while end-to-end systems directly process speech inputs to generate responses without intermediate ASR steps. Each LLM answers times per quesion and use the majority vote as the final answer Wang et al. (2023). QA accuracy is then measured via an exact match and higher accuracy indicates a stronger capability to contextualize and apply spoken content to real-world tasks.
@@ 3.4 Final SIQ Score
Inspired by standard Raven’s progressive matrices John and Raven (2003), a popular human IQ mechanism, we have three steps to compute SIQ: (1) harder samples have higher scores; (2) global standardization making scores among levels computable; (3) dynamic weight among levels in computing final SIQ. For each LLM model and each speech sample , we first compute raw scores for the three dimensions as (taking the negative in subsequent computation), , and .
@@
As one additional study, we also perform normalization based on model scale (i.e., analogous to age factors in human IQ). However, due to the involvement of multiple variables in neural scaling laws Kaplan et al. (2020), we report the non-normalized SIQ and provide preference rankings from a human study to validate its correlation.
@@ 4.1 Datasets
To comprehensively evaluate LLM intelligence across varied domains and real-world challenges, we carefully select two datasets earning22 Rio et al. (2022) and voxpopuli Wang et al. (2021) from the popular OpenASR Leaderboard, while an additional dataset Med-ASR-EN222https://huggingface.co/datasets/jarvisx17/Medical-ASR-EN is chosen from the medical domain, featuring diverse accents and environments. This ensures that our evaluation tests LLM performance in different real-world scenarios, including domain-specific speech and challenging acoustic conditions. Considering the high computational cost of leveraging GPT-4o APIs for QA evaluation and running end-to-end LLM (e.g., Gemini), we extract a subset constructed by the longest audios from each dataset shown in Table 1.
@@ 4.2 Models
We select GPT-4o (OpenAI et al., 2024) as the GER module and use Whisper-large-v2 to generate 5 hypotheses for each audio based on beam search.
@@
For speech-to-text multimodal models, we compare Salmonn (Tang et al., 2024), Qwen2-audio (Chu et al., 2024), desta2 (Lu et al., 2025) and the latest Qwen-2.5-Omni(Xu et al., 2025), we also involve any-to-any multimodel models with speech-to-text capabilities: AnyGPT (Zhan et al., 2024), Baichuan-omni-1.5 (Li et al., 2025), and Gemini-1.5 (Team et al., 2024). We specifically denote Baichuan-omni-1.5 and Gemini-1.5 as “end-to-end-large” since they are either scaled up in training data or model size.
In the understanding level test, we select LLaMA-3.1-8B-Instruct to generate hidden states as responses since related work (Jiang et al., 2024; Liu et al., 2024) has proved the strong embedding capability of LLaMA-based models.
Meanwhile, for the LLM in both ASR + LLM and ASR + GER + LLM to answer QA, we use Qwen2-7b (Yang et al., 2024a) for fair comparisons in size. Preliminary experiments to validate Qwen2-7b and other experiment details are in Appendix B.

### v2

# SpeechIQ: Speech-Agentic Intelligence Quotient Across Cognitive Levels in Voice Understanding by Large Language Models

@@ Abstract
SpeechIQ: Speech-Agentic Intelligence Quotient Across Cognitive Levels in Voice Understanding by Large Language Models
@@ 1 Introduction
The rapid rise of voice understanding Large Language Models (LLM) that can process voice input via the speech portal has ushered in a new paradigm of human-machine interaction reddy1988foundations, where LLM process spoken instructions, infer semantic intent, and execute downstream tasks jelinek1991dynamic; jurafsky1995using; zue2000conversational; tur2005combining; mesnil2014using; kawahara2019spoken; huang2024dynamic; Mahmood_2025. LLM bridge speech and language intelligence sparks1996role, enabling applications from voice assistants to interactive robots. A fundamental prerequisite for deploying robust LLM lies in establishing reliable evaluation metrics for voice understanding. These metrics ensure that the system accurately comprehends voice inputs. Automatic speech recognition (ASR) is the central component of existing cascaded111We refer to two-pass based neural ASR-LM systems hoffmeister2016language; sainath19_interspeech; huang20f_interspeech, which have been widely adopted in the current voice assistants, including Siri, Alexa, and Google Assistant. LLM systems, transcribing voice into text for subsequent LLM processing. Consequently, the research community primarily assesses voice understanding using transcription error metrics hunt1990figures; klakow2002testing, with Word Error Rate (WER) standing as the de facto standard.
@@
However, WER primarily measures the lexical recall of ASR, its transcribed quality does not fully capture the semantic understanding he2011word; chen2018spoken; szymanski2020we and task completion van2020assessing capabilities of LLMs via their speech portal. For instance, WER treats all lexical errors equally, ignoring how transcription inaccuracies propagate to higher-level comprehension (i.e., semantic understanding) and executions (i.e., downstream tasks).
@@
Meanwhile, a growing trend in LLM research is the shift toward end-to-end multi-modal models rubenstein2023audiopalm; radhakrishnan2023whispering; maiti2024voxtlm; hu2024wavllmrobustadaptivespeech; defossez2024moshi; chen2024salm, which bypass explicit ASR transcription and directly map audio inputs to task-specific outputs (e.g., answers, actions). While this paradigm simplifies pipelines and reduces error cascades, its end-to-end voice service obviates the need for generating an intermediate transcript, rendering traditional WER-based evaluation obsolete kuo20_interspeech; tuske21_interspeech. This creates a critical methodological gap: without a reliable and potentially unified metric, comparing voice understanding across architectural choices, such as the modularity of “cascaded neural systems” and the efficiency of “end-to-end mutlimodal models,” remains insufficiently characterized.
@@
The above two limitations mirror a fundamental challenge in assessing artificial intelligence: as Figure 1, human-like intelligence is inherently hierarchical. Drawing on Bloom’s Taxonomy bloom1969taxonomy; 10.1145/1553374.1553380, a cognitive framework where skills ascend from basic remembering to advanced creating, we posit that LLM should be evaluated through analogous levels. In this framework, WER represents the lowest level by measuring verbatim recall while neglecting higher-order abilities such as semantic understanding and task-solving. For instance, two transcripts with identical WER may differ drastically in meaning szymanski2023aren, leading to divergent LLM responses or failed instructions.
@@
To address this gap, we propose a hierarchical IQ test for LLM that evaluates intelligence across three fundamental levels (refer to limitations) aligned with Bloom’s Cognitive Taxonomy: (1) Remember in verbatim: Lexical accuracy is quantified using WER; (2) Understand: Semantic consistency is assessed by comparing LLM responses to ASR outputs and ground-truth transcripts. LLMs are prompted to infer the background context (e.g., domain, speaker intent) and generate a summary of the speech, then measure hidden states similarity between ASR-derived and ground truth-derived responses; (3) Apply: Task-solving capability is tested via multi-choice question-answering (QA) pairs constructed from ground-truth transcripts. LLM answers questions based on speech inputs, with accuracy reflecting real-world utility. Then, we draw inspiration from Raven’s Progressive Matrices raven1998raven; John2003 (i.e., a human IQ test) that aggregates performance across various dimensions into a single score.
@@
Cross-model QA consistency exposes ground-truth errors. By isolating questions most LLM, cannot solve, we create an “unanswerable” set that helps detect hallucinations frieske2024hallucinationsneuralautomaticspeech; Koenecke2024CarelessWS and refine benchmark annotations.
@@ 2.1 ASR Evaluation
ASR systems have long been evaluated using lexical similarity metrics such as WER, which quantifies the Levenshtein distance 10.1145/375360.375365 between ASR outputs and reference transcripts. Metrics like character error rate mackenzie2002character (CER), sentence error rate juffs1996garden (SER), and translation error rate (TER) extend this framework to the sentence or translation level but share WER’s core limitation: treating all errors equally, regardless of semantic impact. While effective for benchmarking transcription fidelity, these metrics do not fully capture how errors propagate to downstream tasks.
Recent work mitigates WER’s semantic insensitivity by incorporating sentence embeddings into evaluation. Sentence similarity-based metrics kim2021semanticdistancenewmetric; kim2022evaluatinguserperceptionspeech and BERTScore zhang2020bertscoreevaluatingtextgeneration leverage pre-trained language models to compute semantic correspondence between ASR hypotheses and ground-truth references. Others propose task-specific metrics, such as Medconcept WER adedeji2024soundhealthcareimprovingmedical which arranges more weights on medical entities in WER computation or severity score whetten-kennington-2023-evaluating that involves sentiment similarity in the evaluation. However, these methods are either confined to specific tasks or require curated labeled data, limiting their generalizability to open-domain LLM.
Efforts to develop hybrid metrics that combine error rate and semantic similarity, such as H_eval sasindran2023hevalnewhybridevaluation, and Sema Sasindran_2024, further enrich the evaluation landscape. Concurrently, machine translation inspired metrics like BLEU Score papineni-etal-2002-bleu and COMET rei-etal-2020-comet, have started to be adapted for ASR to assess fluency and pragmatic adequacy. However, these approaches still emphasize text-to-text alignment between ASR outputs and references, rather than from a view of influencing the responses and actions of downstream LLMs, motivating the proposed SpeechIQ as a multi-dimensional metric.
@@ 2.2 LLM Understanding Systems
In general, speech-based LLMs encompass both the understanding of voice signals and the generation agostinelli2023musiclm; borsos2023audiolm; pmlr-v235-yang24x of vocal or general audio outputs. In this work, we focus exclusively on voice understanding tasks with text outputs as one first step to examine voice intelligence sparks1996role or its potential voice world model ha2018recurrent; matsuo2022deep via LLM backbone, which we denote as LLM. LLM understanding systems are designed to interact shah2018building; lopez2018alexa with users and execute task-oriented instructions via speech-based inputs. This form of “voice-to-text” architecture can be broadly categorized into three main types, all of which achieve state-of-the-art ASR quality on public leaderboards radford2022robustspeechrecognitionlargescale; Chen2023HyPoradiseAO; puvvada2024moreaccuratespeechrecognition:
@@
This approach follows a traditional pipeline-based structure, where an ASR model watanabe2018espnetendtoendspeechprocessing; gulati20_interspeech; radford2022robustspeechrecognitionlargescale transcribes the spoken input, and the 1-best transcription is passed to textual LLM as one operation system wu2024just; dighe2024leveraging for direct response generation.
@@
In this approach, the ASR model first generates multiple possible transcriptions (i.e., hypotheses) for a given speech input, where “texts” are then treated as input features for a many-to-one GER post-editing module Yang_2023; velikovich2024spelling; hori2025delayed. The GER module refines these hypotheses to determine the most accurate transcription, which is subsequently processed by the LLM.
@@
These models are inherently designed to support both speech and text modalities rubenstein2023audiopalm; hu2024wavllmrobustadaptivespeech; lyu2023macawllmmultimodallanguagemodeling; zhang2024speechgptgenscalingchainofinformationspeech. Unlike cascaded approaches, they process audio inputs directly and generate textual outputs without requiring intermediate ASR transcriptions. This paradigm leverages a multimodal learning framework (i.e., a speech embedding encoder radhakrishnan2023whispering or tokenizer borsos2023soundstorm; zhang2024speechtokenizer) to achieve end-to-end speech understanding and reasoning.
In summary, we carefully selected and examined aforementioned three ASR-LLM systems. System 1 serves as the most cost-effective solution 196553 for voice interactions with LLMs. System 2 is designed to decode texts that offer richer information chan-etal-2023-ic3; yang2024large for textual applications. System 3 highlights a low-latency, multimodal embedding injection approach for large language models openai2024gpt4ocard; geminiteam2024gemini15unlockingmultimodal. Next, we introduce our proposed method, which corresponds to Bloom’s cognitive taxonomy.
@@ 3.2 Understand: Similarity Metric
To quantify the impact of ASR errors on semantic comprehension, we measure the deviation in LLM responses caused by ASR transcription errors. Specifically, inspired by recent work (jiang-etal-2024-scaling; inproceedings) showing that instructing LLMs to generate one word is a simple and straightforward approach to representing semantic meanings, we evaluate LLM-generated responses via two key questions: (1) : The background scenario of the speech [] in one word is. This helps resolve ambiguities and measure whether the ASR content preserves core contextual meaning; (2) : The summary of this speech [] in one word is. This assesses the extent to which ASR errors affect the LLM’s high-level understanding. Instead of directly comparing LLM-generated words, we use the last layer of hidden states of LLM for generating the next token as embeddings and compute their cosine similarity with the embeddings generated from the ground truth transcription:
@@ 3.3 Apply: QA Accuracy
During QA generation, we leverage GPT-4o and prompt it to focus on either the core concept or information details in the transcription. (Appendix A) To mitigate potential errors in generated QA, we also employ GPT-4o itself and Gemini-1.5-flash to answer these questions using the ground-truth transcription. Questions, where GPT-4o or Gemini-1.5-flash fails to produce correct answers, are discarded and regenerated. Then during evaluation, cascaded systems answer questions based on ASR transcriptions, while end-to-end systems directly process speech inputs to generate responses without intermediate ASR steps. Each LLM answers times per quesion and use the majority vote as the final answer wang2023selfconsistencyimproveschainthought. QA accuracy is then measured via an exact match and higher accuracy indicates a stronger capability to contextualize and apply spoken content to real-world tasks.
@@ 3.4 Final SIQ Score
Inspired by standard Raven’s progressive matrices John2003, a popular human IQ mechanism, we have three steps to compute SIQ: (1) harder samples have higher scores; (2) global standardization making scores among levels computable; (3) dynamic weight among levels in computing final SIQ. For each LLM model and each speech sample , we first compute raw scores for the three dimensions as (taking the negative in subsequent computation), , and .
@@
As one additional study, we also perform normalization based on model scale (i.e., analogous to age factors in human IQ). However, due to the involvement of multiple variables in neural scaling laws Kaplan2020ScalingLF, we report the non-normalized SIQ and provide preference rankings from a human study to validate its correlation.
@@ 4.1 Datasets
To comprehensively evaluate LLM intelligence across varied domains and real-world challenges, we carefully select two datasets earning22 delrio2022earnings22practicalbenchmarkaccents and voxpopuli wang-etal-2021-voxpopuli from the popular OpenASR Leaderboard, while an additional dataset Med-ASR-EN222https://huggingface.co/datasets/jarvisx17/Medical-ASR-EN is chosen from the medical domain, featuring diverse accents and environments. This ensures that our evaluation tests LLM performance in different real-world scenarios, including domain-specific speech and challenging acoustic conditions. Considering the high computational cost of leveraging GPT-4o APIs for QA evaluation and running end-to-end LLM (e.g., Gemini), we extract a subset constructed by the longest audios from each dataset shown in Table 1.
@@ 4.2 Models
We compare the following widely used ASR models: Whisper-Large-v2, Whisper-Large-v3 (radford2022robustspeechrecognitionlargescale), Canary (puvvada2024moreaccuratespeechrecognition), and ESPnet (owsm_ctc_v3.1_1B) (watanabe2018espnetendtoendspeechprocessing)
@@
We select GPT-4o (openai2024gpt4ocard) as the GER module and use Whisper-large-v2 to generate 5 hypotheses for each audio based on beam search.
@@
For speech-to-text multimodal models, we compare Salmonn (tang2024salmonngenerichearingabilities), Qwen2-audio (chu2024qwen2audiotechnicalreport), desta2 (lu2025desta2developinginstructionfollowingspeech) and the latest Qwen-2.5-Omni(xu2025qwen25omnitechnicalreport), we also involve any-to-any multimodel models with speech-to-text capabilities: AnyGPT (zhan2024anygptunifiedmultimodalllm), Baichuan-omni-1.5 (li2025baichuanomni15technicalreport), and Gemini-1.5 (geminiteam2024gemini15unlockingmultimodal). We specifically denote Baichuan-omni-1.5 and Gemini-1.5 as “end-to-end-large” since they are either scaled up in training data or model size.
In the understanding level test, we select LLaMA-3.1-8B-Instruct to generate hidden states as responses since related work (jiang-etal-2024-scaling; inproceedings) has proved the strong embedding capability of LLaMA-based models.
Meanwhile, for the LLM in both ASR + LLM and ASR + GER + LLM to answer QA, we use Qwen2-7b (yang2024qwen2technicalreport) for fair comparisons in size. Preliminary experiments to validate Qwen2-7b and other experiment details are in Appendix B.